divide 4 by 8/7 I can't remember how to do this thanks
Answer by fractalier(6550)

To divide 4 by 8/7, you turn it into a multiplication problem and invert the 8/7...so we get
4 / (8/7) = 4 * (7/8) = 28/8 = 7/2 or 3 1/2